# Optical systems with translational invariance

### v2

@@ 3.4.2 Rest of the terms
$B_{01}$ $=$ $\left(\alpha_{c2}+\alpha_{e3}L_{x}\right)L_{y}^{2},\quad B_{02}=\left(\alpha_{a2}+2\alpha_{c2}L_{x}+\alpha_{e3}L_{x}^{2}+\alpha_{e2}L_{y}^{2}\right)L_{y}$ (105)
$B_{11}$ $=$ $\frac{\alpha_{e3}}{2}L_{y}^{2},\quad B_{12}=\frac{3\alpha_{e2}}{2}L_{y}^{2},\quad B_{13}=2\left(\alpha_{c2}+\alpha_{e3}L_{x}\right)L_{y},$ (106)
$B_{22}$ $=$ $\alpha_{e2}L_{y},\quad B_{23}=\alpha_{e3}L_{y}.$ (107)
@@
$L_{y}=W\chi\left(1+\frac{\chi^{2}}{3}\right)+\mathcal{O}(\chi^{5}),$ (108)
where $\chi\approx\eta_{0}$ and the comment below Eq. (103) is also applicable. The final expressions for $B_{ij}$ are shown in Table 3. The upper and lower signs in this table and in Eq. (108) correspond to the entrance pupil located in front of the surface and behind it, respectively.
Table 3: Aberration coefficients $B$ for a cylindrical surface in case of displaced stop
$B_{ij}$ Refracting surface of Eq. (38) Reflecting surface of Eq. (80)
$B_{22}$  $-\frac{n_{0}(W\chi)}{2s_{0}^{3}}\left[1-\frac{s_{0}^{2}}{s_{1}^{2}}\,\frac{(1+\beta)^{4}-\beta^{4}}{\beta}\right]$  $-\frac{n_{0}(W\chi)}{2s_{0}^{3}}\,\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}}$
$B_{23}$  $-\frac{n_{0}(W\chi)}{2s_{0}^{2}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\Gamma_{R}$  $-\frac{n_{0}(W\chi)}{s_{0}^{2}R}\,\frac{\mu+1}{\mu-1}$
$B_{11}$  $-\frac{n_{0}(W\chi)^{2}}{4s_{0}^{2}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\Gamma_{R}$  $-\frac{n_{0}(W\chi)^{2}}{2s_{0}^{2}R}\,\frac{\mu+1}{\mu-1}$
$B_{12}$  $-\frac{3n_{0}(W\chi)^{2}}{4s_{0}^{3}}\left[1-\frac{s_{0}^{2}}{s_{1}^{2}}\,\frac{(1+\beta)^{4}-\beta^{4}}{\beta}\right]$  $-\frac{3n_{0}(W\chi)^{2}}{4s_{0}^{3}}\,\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}}$
$B_{13}$  $-\frac{n_{0}W^{2}}{s_{0}^{2}}\psi\chi\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\left(\frac{1}{W}\pm\frac{1}{R}\right)$  $\mp\frac{2n_{0}}{s_{0}^{2}R}W^{2}\psi\chi\,\frac{\mu+1}{\mu-1}$
$B_{01}$  $-\frac{n_{0}}{2s_{0}^{2}}W^{3}\psi\chi^{2}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\left(\frac{1}{W}\pm\frac{1}{R}\right)$  $\mp\frac{n_{0}}{s_{0}^{2}R}W^{3}\psi\chi^{2}\frac{\mu+1}{\mu-1}$
$B_{02}$  $-\frac{n_{1}-n_{0}}{R}(W\chi)\left(1+\frac{\chi^{2}}{3}\right)-\frac{n_{0}(W\chi)^{3}}{2s_{0}^{3}}\left[1-\frac{s_{0}^{2}}{s_{1}^{2}}\,\frac{(1+\beta)^{4}-\beta^{4}}{\beta}\right]$  $-n_{0}\frac{(W\chi)}{s_{0}}\left(1+\frac{\chi^{2}}{3}\right)\frac{\mu-1}{\mu}-n_{0}\frac{W^{3}}{s_{0}^{2}R}\psi^{2}\chi\frac{\mu+1}{\mu-1}$
 $-\frac{n_{0}W^{3}}{2s_{0}^{2}}\psi^{2}\chi\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\left[2\frac{s_{0}}{W}\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)+\Gamma_{R}\right]$  $-\frac{n_{0}(W\chi)^{3}}{2s_{0}^{3}}\,\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}}$
@@ 4.1 Generic discussion
where $f$ refers to the final surface. At the fourth order, each $\Phi_{i}^{(4)}$ can be replaced by the sum of the expressions given in Eqs. (92) and (3.4.2), with the appropriate $x^{\prime}$ and $y^{\prime}$ at each surface denoted by $x_{i}$ and $y_{i}$, respectively. Correspondingly, we denote the refraction indices $n_{0i}$ and $n_{1i}$, the distance $W_{i}$ from the $i$th surface to its entrance pupil, the angles $\psi_{i}$, $\chi_{i}$, etc.
@@ 4.1.1 $XZ$ plane
The Eq. (112) allows to calculate the transverse aberration with the marginal ray height at the last surface. This aberration can also be expressed in terms of the marginal ray height at some other surface (e.g., at the OS entrance pupil), say the $l$th one, as follows
@@ 4.1.2 $YZ$ plane
which also can be rewritten in a form similar to Eq. (114).
@@ 4.2 Cylindrical optical system of Cassegrain type
$A_{011}(L_{x1}=0)$ $=$ $\alpha_{d}=0,\quad A_{111}(L_{x1}=0)=\frac{\alpha_{a1}}{2}=-\frac{\theta_{0}^{2}}{2R_{1}},$ (120)
@@
From the law of reflection with the replacement discussed below Eq. (103) we find
@@
$A_{012}$ $=$ $\frac{(W\theta_{0})^{3}}{2R_{2}^{3}}\left[\kappa_{2}-\left(\frac{\mu+1}{\mu-1}\right)^{2}+3R_{2}^{2}\left(\frac{1}{W}-\frac{1}{R_{2}}\right)^{2}-R_{2}^{2}\left(\frac{1}{W}-\frac{1}{s_{02}}\right)^{2}\right],$ (123)
@@
where $W\equiv W_{2}$, $\mu\equiv\mu_{2}$ and the overall signs are taken positive since the mirror is convex. From Eq. (115) for the marginal ray height at the primary $x_{l}=x_{1}$ we have
@@
For the rays in the $YZ$ plane ($L_{xi}=0$) for the primary ($L_{y1}=0$) we have only two nonzero coefficients:
$\tilde{B}_{11}$ $=$ $\frac{\alpha_{a2}}{2}=-\frac{1}{2}\left(\frac{1}{s_{01}}+\frac{1}{s_{11}}\right)=-\frac{1}{2f_{1}},$ (132)
$\tilde{B}_{31}$ $=$ $\frac{\alpha_{e2}}{2}=-\frac{1-4\mu_{1}+6\mu_{1}^{2}-3\mu_{1}^{3}}{8s_{01}^{3}\mu_{1}^{3}}=\frac{1}{8f_{1}^{3}}.$ (133)
From the law of reflection in Eq. (87) and the agreement on $\chi$ below Eq. (108) we have
$\chi_{2}=\pi-\eta_{11}=\eta_{01}\equiv\eta_{0}.$ (134)
Then for the convex secondary we obtain
$\tilde{B}_{02}$ $=$ $B_{022}(L_{x2}=0)=-\frac{(W\eta_{0})}{s_{02}}\frac{1-\mu}{\mu}+\mathcal{O}(\theta_{0}^{2}\eta_{0},\eta_{0}^{3}),$ (135)
$\tilde{B}_{12}$ $=$ $A_{122}(L_{x2}=0)+B_{122}=-\,\frac{1}{2s_{02}}\frac{1-\mu}{\mu}+\frac{3(W\eta_{0})^{2}}{2s_{02}^{3}}\,\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}},$ (136)
$\tilde{B}_{22}$ $=$ $B_{222}=\frac{(W\eta_{0})}{2s_{02}^{3}}\,\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}},$ (137)
$\tilde{B}_{32}$ $=$ $A_{322}=\frac{1}{8s_{02}^{3}}\,\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{\mu^{3}}.$ (138)
From
$\tilde{B}_{js}^{(1)}=\tilde{B}_{j1}+\tilde{B}_{j2}\left(\frac{y_{2}}{y_{1}}\right)^{j+1},$ (139)
for $y_{1}=y_{2}$, using the relation of $s_{02}=-k_{x}f_{1}$ (for the virtual object for the secondary), we have
$\tilde{B}_{3s}^{(1)}=\frac{1}{8f_{1}^{3}}F_{YZ}(\mu,k_{x}),$ (140)
where we took $y_{1}=y_{2}$, and the new function is defined as
$F_{YZ}(\mu,k_{x})=1-\frac{1-4\mu+6\mu^{2}-3\mu^{3}}{(\mu k_{x})^{3}}.$ (141)
Finally, using $s_{12}=-\mu s_{02}=\mu k_{x}f_{1}$, the transverse elliptic aberration for the $YZ$ plane is given by
${\rm TA}_{3YZ}=\frac{\mu k_{x}}{2}\,\frac{y_{1}^{3}}{f_{1}^{2}}F_{YZ}(\mu,k_{x}).$ (142)
Expressions for the rest of the transverse aberrations for the rays in the $XZ$ and $YZ$ planes can be determined using the same procedure. Condition $F_{XZ}=0$ for zero elliptic aberration in the $XZ$ plane coincides with the known condition of the spherical aberration compensation for the rotationally symmetric Cassegrain telescope. One of the possibilities for such compensation provides the classical Cassegrain telescope with the paraboloidal primary and hyperboloidal secondary, i.e.:
@@
Condition $F_{YZ}=0$ for vanishing elliptic aberration in the $YZ$ plane results in the equation
$(3+k_{x}^{3})\mu^{3}-6\mu^{2}+4\mu-1=0,$ (144)
which is satisfied for close to unity values of $\mu$ for reasonable values of $k_{x}$, but does not make much sense for the isolated discussed cylindrical OS. However, if this OS is embedded to an anamorphic OS, which optical axis is the $Z$ axis, then the functions $F_{XZ}$ and $F_{YZ}$ or their generalizations enter the equations for the full system leading aberrations compensation.
@@ 5 Conclusion
In this paper the method of calculation of monochromatic aberrations for the optical systems with the symmetry of a generic cylinder have been developed. In particular, the aberration coefficients for calculations of the aberrations for an arbitrary paraxial ray were obtained. Interestingly, the second coma is vanishing together with the distortion for a single mirror surface ($\alpha_{c2}=0=\alpha_{d}$ in Table 1). Further, the derived analytical results have been applied to an example optical system, namely, the cylindrical analog of the Cassegrain telescope.

### v3

@@ 3.4.2 Rest of the terms
$B_{01}$ $=\left(\alpha_{c2}+\alpha_{e3}L_{x}\right)L_{y}^{2},\quad B_{02}=\left(2\alpha_{c2}+\alpha_{e3}L_{x}\right)L_{x}L_{y},$ (101)
$B_{11}$ $=\frac{\alpha_{e3}}{2}L_{y}^{2},\quad B_{12}=0,\quad B_{13}=2\left(\alpha_{c2}+\alpha_{e3}L_{x}\right)L_{y},$ (102)
$B_{22}$ $=0,\quad B_{23}=\alpha_{e3}L_{y}.$ (103)
@@
$L_{y}=W\tan\chi=W\chi\left(1+\frac{\chi^{2}}{3}\right)+\mathcal{O}(\chi^{5}),$ (104)
where $\chi\approx\eta_{0}$ and the comment below (99) is also applicable. The final expressions for $B_{ij}$ are shown in Tables 4 and 5. The upper and lower signs in these tables and in (104) correspond to the entrance pupil located in front of the surface and behind it, respectively.
Table 4: Aberration coefficients $B$ for the refracting cylindrical surface of (38) in case of displaced stop
$B_{23}$  $-\frac{n_{0}(W\chi)}{2s_{0}^{2}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\Gamma_{R}$
$B_{11}$  $-\frac{n_{0}(W\chi)^{2}}{4s_{0}^{2}}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\Gamma_{R}$
$B_{13}$  $-\frac{n_{0}W^{2}}{s_{0}^{2}}\psi\chi\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\left(\frac{1}{W}\pm\frac{1}{R}\right)$
$B_{01}$  $-\frac{n_{0}}{2s_{0}^{2}}W^{3}\psi\chi^{2}\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\left(\frac{1}{W}\pm\frac{1}{R}\right)$
$B_{02}$  $-\frac{n_{0}W^{3}}{2s_{0}^{2}}\psi^{2}\chi\left(1-\frac{n_{0}^{2}}{n_{1}^{2}}\right)\left[\pm 2\left(\frac{1}{W}\mp\frac{1}{s_{0}}\right)+\Gamma_{R}\right]$
@@ 4.1 Generic discussion
where $f$ refers to the final surface. At the fourth order, each $\Phi_{i}^{(4)}$ can be replaced by the sum of the expressions given in Eqs. (3.4.1) and (3.4.2), with the appropriate $x^{\prime}$ and $y^{\prime}$ at each surface denoted by $x_{i}$ and $y_{i}$, respectively. Correspondingly, we denote the refractive indices $n_{0i}$ and $n_{1i}$, the distance $W_{i}$ from the $i$th surface to its entrance pupil, the angles $\psi_{i}$, $\chi_{i}$, etc.
Table 5: Aberration coefficients $B$ for the reflecting cylindrical surface of (77) in case of displaced stop
$B_{23}$  $-\frac{n_{0}(W\chi)}{s_{0}^{2}R}\,\frac{\mu+1}{\mu-1}$
$B_{11}$  $-\frac{n_{0}(W\chi)^{2}}{2s_{0}^{2}R}\,\frac{\mu+1}{\mu-1}$
$B_{13}$  $\mp\frac{2n_{0}}{s_{0}^{2}R}W^{2}\psi\chi\,\frac{\mu+1}{\mu-1}$
$B_{01}$  $\mp\frac{n_{0}}{s_{0}^{2}R}W^{3}\psi\chi^{2}\frac{\mu+1}{\mu-1}$
$B_{02}$  $-n_{0}\frac{W^{3}}{s_{0}^{2}R}\psi^{2}\chi\frac{\mu+1}{\mu-1}$
@@ 4.1.1 $XZ$ plane
The (108) allows to calculate the transverse aberration with the marginal ray height at the last surface. This aberration can also be expressed in terms of the marginal ray height at some other surface (e.g., at the OS entrance pupil), say the $l$th one, as follows
@@ 4.1.2 $YZ$ plane
which also can be rewritten in a form similar to (110). However, the numerical values of all $\tilde{B}$ in (113) are vanishing.
@@ 4.2 Cylindrical optical system of Cassegrain type
$A_{011}(L_{x1}=0)$ $=\alpha_{d}=0,\ \ A_{111}(L_{x1}=0)=\frac{\alpha_{a1}}{2}=-\frac{\theta_{0}^{2}}{2R_{1}},$ (116)
@@
From the law of reflection with the replacement discussed below (99) we find
@@
$A_{012}=\frac{(W\theta_{0})^{3}}{2R_{2}^{3}}$
×$\times\left[\kappa_{2}-\left(\frac{\mu+1}{\mu-1}\right)^{2}+3R_{2}^{2}\left(\frac{1}{W}-\frac{1}{R_{2}}\right)^{2}-R_{2}^{2}\left(\frac{1}{W}-\frac{1}{s_{02}}\right)^{2}\right],$ (119)
@@
where $W\equiv W_{2}$, $\mu\equiv\mu_{2}$ and the overall signs are taken positive since the mirror is convex. From (111) for the marginal ray height at the primary $x_{l}=x_{1}$ we have
@@
Expressions for the rest of the transverse aberrations for the rays in the $XZ$ and $YZ$ planes can be determined using the discussed procedure. Condition $F_{XZ}=0$ for zero elliptic aberration in the $XZ$ plane coincides with the known condition of the spherical aberration compensation for the rotationally symmetric Cassegrain telescope. One of the possibilities for such compensation provides the classical Cassegrain telescope with the paraboloidal primary and hyperboloidal secondary, i.e.:
@@
If the discussed OS is embedded to an anamorphic OS, which optical axis is the $Z$ axis, then the function $F_{XZ}$ enters respective equation for the full system leading aberrations compensation.
@@ 5 Conclusion
In this paper the method of calculation of monochromatic aberrations for the optical systems with the symmetry of a generic cylinder have been developed. In particular, the aberration coefficients for calculations of the aberrations for an arbitrary paraxial ray were obtained in certain approximations. Interestingly, the second coma is vanishing together with the distortion for a single mirror surface ($\alpha_{c2}=0=\alpha_{d}$ in Table 1). Further, the derived analytical results have been applied to an example optical system, namely, the cylindrical analog of the Cassegrain telescope.
We notice that choice of suitable approximations, in general, depends on the parameters of the optical system and requires further investigation.